(2,4) (x,8) slope = -2
What is x? And how do you find x?

Answer by richwmiller(17219)

(4-8)/(2-x)=-2
-4=-2*2+2x
-4=-4+2x
add 4
0=2x
x=0
check
4-8/2-0=-4/2=-2